OpenStudy (anonymous):

Find the equation of the line with slope that is -3 and passing through the point (2,-1)

OpenStudy (anonymous):

hey heres a good video that will help you on how to do alot of these videos > http://www.youtube.com/watch?v=OjEKoXBZf7w

OpenStudy (anonymous):

the line be y=mx +c since it passes through (2,-1) -1=2*m+c since slope is -3, m=-3 -1=2*-3+c -1=-6+c c=5 the line is hence y=-3x+5 3x+y-5=0